OpenStudy (anonymous):

A Triangle has a leg that measures 3 and a hypotenuse of 9. Which leg measure would form a right triangle? A)12 B)90 C)✔️72 D)✔️90

OpenStudy (solomonzelman):

`3² + b² = c² `

OpenStudy (solomonzelman):

Solve for b (only take the positive value of the square root that you get)

OpenStudy (solomonzelman):

I meant `3² + b² = 9² `

OpenStudy (solomonzelman):

This will take another shape, (rearrangement) b=√(3²+9²)=√(9+81)=√90=√(9 × 10)= 3√10 ≈ 9.487